OpenStudy (anonymous):

what is te exact value of thr expression sqrt72 - sqrt 8 + sqrt 128

OpenStudy (anonymous):

72 = 2 * 36 8 = 2^3 128 = 2^7 Can you continue?

OpenStudy (anonymous):

Ii dont know how to do itt

OpenStudy (anonymous):

36 = 6² √( 2 * 6² ) = 6 √ 2 Do you understand this step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you continue for √ 2³ = ....?

OpenStudy (anonymous):

i told you, i dont get this! lol

OpenStudy (anonymous):

Why did you say "Yes" ??? Yes means what???